OpenStudy (anonymous):

In triangle ABC, AB=2x, AC=x, BC=21 and angle BAC=120. Calculate the value of x. I think i have to use cosine rule but dont get it will fan and medal

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

law of cosines

OpenStudy (anonymous):

i know so what i did was 21^2= x^2+4x^2-(2 times 2x times x cos 120)? is this right?

OpenStudy (anonymous):

draw the triangle, it will be easier to see and solve

OpenStudy (anonymous):



OpenStudy (anonymous):

so \[21^2=x^2+\left( 2x \right)^2-2\left( x \right)\left( 2x \right)\cos120^\circ\]

OpenStudy (anonymous):

yah i drew it yah

OpenStudy (anonymous):

from this you should get a quadratic in x. solve for x

OpenStudy (anonymous):

then is it x^2 cos 120=21^2?

OpenStudy (anonymous):

\(441=5x^2-4x^2(-1/2)\) yeah?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

what does this give me?

OpenStudy (anonymous):

you tell me... work it out

OpenStudy (anonymous):

ah 7x^2=21^2? right

OpenStudy (anonymous):

so x would be 7.94...

OpenStudy (anonymous):

so \(x^2 = 63 \Rightarrow x = \sqrt{63}\)

OpenStudy (anonymous):

good to go?

OpenStudy (anonymous):

yup